# Recall, Robustness, and Lexicographic Evaluation

### v5

@@ Abstract.
Although originally developed to evaluate sets of items, recall is often used to evaluate rankings of items, including those produced by recommender, retrieval, and other machine learning systems. The application of recall without a formal evaluative motivation has led to criticism of recall as a vague or inappropriate measure. In light of this debate, we reflect on the measurement of recall in rankings from a formal perspective. Our analysis is composed of three tenets: recall, robustness, and lexicographic evaluation. First, we formally define ‘recall-orientation’ as the sensitivity of a metric to a user interested in finding every relevant item. Second, we analyze recall-orientation from the perspective of robustness with respect to possible content consumers and providers, connecting recall to recent conversations about fair ranking. Finally, we extend this conceptual and theoretical treatment of recall by developing a practical preference-based evaluation method based on lexicographic comparison. Through extensive empirical analysis across three recommendation tasks and 17 information retrieval tasks, we establish that our new evaluation method, lexirecall, has convergent validity (i.e., it is correlated with existing recall metrics) and exhibits substantially higher sensitivity in terms of discriminative power and stability in the presence of missing labels. Our conceptual, theoretical, and empirical analysis substantially deepens our understanding of recall and motivates its adoption through connections to robustness and fairness.
@@ 1. Introduction
Researchers use ‘recall’ to evaluate rankings across a variety of retrieval, recommendation (Zhao et al., 2022), and machine learning tasks (Sajjadi et al., 2018; Rudinger et al., 2015; Flach and Kull, 2015). ‘Recall at $k$ items’ ($\text{R}_{k}$) and R-Precision (RP) are popular metrics used for measuring recall in rankings. And, while there is a colloquial interpretation of recall as measuring coverage (as it might be rightfully interpreted in set retrieval), the research community is far from a principled understanding of recall metrics for rankings. Nevertheless, authors continue to informally refer to evaluation metrics as more or less ‘recall-oriented’ or ‘precision-oriented’ without a formal definition of what this means or quantifying how existing metrics relate to these constructs (Kazai and Lalmas, 2006; Montazeralghaem et al., 2020; Dai et al., 2017; Mohammad et al., 2018; Li et al., 2014; Diaz, 2015).
@@
The lack of a principled understanding of or motivation for recall has caused some to question whether recall is useful as a construct at all. Cooper (1968) argues that recall-orientation is inappropriate because user search satisfaction depends on the number of items the user is looking for, which may be fewer than all of the relevant items. Zobel et al. (2009) refute several informal justifications for recall: persistence (the depth a user is willing to browse), cardinality (the number of relevant items found), coverage (the number of user intents covered), density (the rank-locality of relevant items), and totality (the retrieval of all relevant items). So, while many ranking experiments compute recall metrics, precisely what and why we are measuring remains unclear.
@@
In this light, we approach the measurement of recall in rankings from a formal perspective, with an objective of proposing a new interpretation of recall with precise conceptual and theoretical grounding. Our analysis is composed of three interrelated tenets: recall, robustness, and lexicographic evaluation. First, we consider recall an essentially contested construct (Gallie, 1955; Jacobs and Wallach, 2021): it is a high level construct believed to be important but with different, conflicting interpretations, as suggested by Zobel et al. (2009). By adopting the interpretation of recall as finding the totality of relevant items, we formally define ‘recall-orientation’ as sensitivity to movement of the bottom-ranked relevant item. Although simple, this definition of recall connects to both early work in position-based evaluation as well as recent work in technology-assisted review. Moreover, by formally defining recall orientation, we can design a new metric, total search efficiency, that precisely measures recall. Second, we consider robustness another essentially contested construct, again with different, conflicting interpretations (Drenkow et al., 2021). By adopting an interpretation of robustness as the effectiveness for the worst-off user, we can connect it to our notion of recall-orientation. We demonstrate that recall—and totality specifically—is aligned with worst-case robustness. Finally, we extend this conceptual and theoretical treatment of recall by developing a practical preference-based evaluation method based on lexicographic comparison. We present a conceptual relationship between recall, robustness, and lexicographic evaluation in Figure 1. Through extensive empirical analysis across 17 TREC tracks and three recommendation tasks, we establish that our new evaluation method, lexicographic recall or lexirecall, is correlated with existing recall metrics but exhibits substantially higher discriminative power and stability in the presence of missing labels. While conceptually and theoretically grounded in notions of robustness and fairness, lexirecall pragmatically is appropriate when we are concerned with understanding system behavior for users who are focused on finding all relevant items in the same way that experiments adopt reciprocal rank as a high-precision metric. This paper, by focusing on worst-case performance, also complements existing work that has focused on average case (Diaz and Ferraro, 2022) and best-case (Diaz, 2023) preference-based evaluation. Our conceptual, theoretical, and empirical analysis substantially deepens our understanding of recall as a construct and motivates its adoption through connections to robustness and fairness constructs.
@@ 2. Preliminaries
We consider ranking systems that are designed to satisfy users [^1] We adopt user as a general term for searcher in the information retrieval context and content consumer in the multi-stakeholder recommendation context. with a specific information need in mind. Although a user’s information need is never directly revealed to the ranking system, the user expresses an observable request to ranking system. A request can be explicit (e.g., a query or question in the context of text-based search) or implicit (e.g., engagement or rating history in the context of recommendation).
@@
The objective of ranking evaluation is to determine the quality of a permutation $\pi\in S_{n}$ for the user. In the remainder of this section, we will detail precisely how we do this.
@@ 2.1. Relevance
The relevance of an item refers to its value with respect to a user’s information need. In this work, we focus on binary relevance, an approach regularly used in information retrieval and recommender system literature, especially when measuring recall. [^2] We discuss how our analysis extends in ordinal grades and preferences in Section 7.2.3.
Let $\mathcal{R}\subset\mathcal{D}$ be the set of items labeled relevant to the request where $m=|\mathcal{R}|$.
@@
Figure 2. Relevance Projection with Imputation. Given set of relevant item ids =$\mathcal{R}=\{083,107,511,668,820,977\}$, relevance projection of an incomplete top-10 ranking $\pi$ and relevant set $\mathcal{R}$ to a $m\times 1$ vector of positions $p$. We also show the inverse projection vector $\overline{p}$ of items at specific recall levels.
@@ 2.2. Permutation Imputation
Pessimistic imputation is a conservative placement of the unretrieved relevant items and is well-aligned with our interest in robust performance. Moreover, it is consistent with behavior of metrics like rank-biased precision, which implicitly applies an exposure of 0 for unretrieved relevant items (i.e., for large $n$, $\lim_{i\rightarrow n}\gamma^{i}\approx 0$); and average precision [^3] As defined in trec_eval., which implicitly applies an exposure of 0 for unretrieved relevant items (i.e., for large $n$, $\lim_{i\rightarrow n}\frac{1}{i}\approx 0$). We show an example of projection with imputation in Figure 2.
@@ 2.3. Measuring Effectiveness
When evaluating an ranking system, we consider two sets of important stakeholders: users and providers. Users approach the system with information needs and requests and ultimately define what is relevant. Providers contribute items to the search system which serve to satisfy users. Each item in the corpus is attributable, explicitly or not, to a content provider. In this section, we will characterize a broad family of evaluation metrics for these two sets of users that will allow us, in subsequent sections, to define formal notions of recall and robustness.
@@ 2.3.1. Measuring Effectiveness for Users
The product $e(p_{i})z(i,m)$ is a decomposition of what Carterette refers to as a ‘discount function’ into an explicit function that models exposure and another that addresses any recall normalization.
Within the set of recall-level metrics, we are further interested in the sub-class of metrics that satisfy the following criteria for ‘top-heaviness’.
Definition 2.2 ().
@@
Top-heaviness indicates that, in the event that there are unjudged, relevant items above the currently highest-ranked relevant items, the metric value must be greater than or equal to the metric value computed over the incomplete judgments. Because we deal with metrics that include functions of $m$, this is not an obvious property, but one that will be important as we consider incomplete judgments and relationships between possible users in Section 4.
@@
Beyond classic ranking metrics, top-heavy recall-level metrics include non-traditional metrics such those based on linear discounting (e.g., $e_{\text{lin}}(i)=1-\frac{i}{n}$). This results in a much broader class of metrics than those normally considered, for example, in the formal analysis of ranking metrics (Moffat, 2013; Amigó et al., 2013; Ferrante et al., 2015). As a result, while all top-heavy recall-level metrics satisfy some formal properties ranking evaluation metrics, large subsets of top-heavy recall-level metrics may satisfy more. A detailed analysis of formal properties of top-heavy recall-level metrics can be found in Appendix A. As mentioned before, we can contrast this with Carterette’s decomposition which focuses on the decomposition of metrics into gain and discount components. In our case, we do not model gain, since we deal with binary relevance. Our exposure and normalization functions, then, precisely define a subset of Carterette’s discount functions that do not fit into his metric taxonomy since they do not consider recall normalization.
@@ 2.3.2. Measuring Effectiveness for Providers
where $m$ and $p$ are based on $\mathcal{R}$. Unless necessary, we will drop the subscript $e$ from $\eta$ for clarity. We summarize our notation in Table 1.
Table 1. Notation
$\mathbb{Z}^{+}$
positive integers
$\mathbb{Z}^{*}$
non-negative integers
$\mathbb{R}^{+}$
positive reals
$\mathbb{R}^{*}$
non-negative reals
$S_{k}$
set of all permutations of $k$ items
$\mathcal{A}^{+}$
non-empty subsets of $\mathcal{A}$
$\text{I}:X\rightarrow\{0,1\}$
indicator function (i.e., returns 1 if $X$ is true; 0 otherwise)
$x_{-i}$
reverse index (i.e., $x_{k-i+1}$ for the $k$-dimensional vector $x$)
$\mathcal{D}$
corpus
$\mathcal{R}$
relevant set
$n$
size of corpus (i.e., $|\mathcal{D}|$)
$m$
size of relevant set (i.e., $|\mathcal{R}|$)
$\tilde{n}$
number of items retrieved
$S_{n}$
set of all permutations of $\mathcal{D}$
$\tilde{S}_{n}$
subset of $S_{n}$ generated by multiple systems for a fixed request
$\pi$
permutation of $\mathcal{D}$
$p$
sorted positions of relevant items
$\overline{p}$
item ids of relevant items sorted by position
$C$
number of unique permutations in $S_{n}$ for a given $p$ (i.e., $m!(n-m)!$)
$\mu:S_{n}\times\mathcal{D}^{+}\rightarrow\mathbb{R}^{*}$
user evaluation metric
$\eta:S_{n}\times\mathcal{D}^{+}\rightarrow\mathbb{R}^{*}$
provider evaluation metric
$\mu(\pi,\mathcal{R})\mathrel{\stackrel{{\scriptstyle\makebox[0.0pt]{\mbox{\tiny rank}}}}{{=}}}\mu^{\prime}(\pi,\mathcal{R})$
$\mu$ and $\mu^{\prime}$ rank $S_{n}$ identically
$e:\mathbb{Z}^{+}\rightarrow\mathbb{R}^{*}$
exposure of position
$z:\mathbb{Z}^{+}\times\mathbb{Z}^{+}\rightarrow\mathbb{R}^{*}$
normalization function
$\mathcal{U}$
set of possible user information needs for a request
$\mathcal{V}$
set of possible providers for a request
@@ 2.4. Evaluation Method Desiderata
    • System sensitivity. Is the evaluation method better able to distinguish between systems compared to existing methods? (Section 6.2.3)
@@ 3. Recall
As mentioned in Section 1, the description of ranked ranking metrics as ‘recall-oriented’ remains poorly defined, leaving the formal analysis of metrics for recall-orientation difficult. From a technical point of view, some work considers recall-orientation to be a binary criteria, dependent on whether a metric includes the recall base (i.e., $\mathcal{R}$) in order to be computed (Kazai and Lalmas, 2006; Sakai, 2006). This would include metrics that compute set-based recall at some rank cutoff (Tomlinson and Hedin, 2017; Cormack and Grossman, 2017) as well as metrics like AP and $\mathrm{NDCG}$. Using a set-based recall metric is particularly well-suited for recall-orientation in early stages of multi-stage ranking (Macdonald et al., 2013; Mohammad et al., 2018). A binary notion of recall-orientation does not capture that some metrics may be more recall-oriented than others. This captured, in part, by references to recall-orientation as related to the depth in the ranking considered by the user (Diaz, 2015). More frequently, authors appeal to metrics like AP and $\text{R}_{1000}$ as being recall-oriented without clear discussion of what this means (Montazeralghaem et al., 2020; Dai et al., 2017; Li et al., 2014; Hashemi et al., 2020). On the other hand, both Mackie et al. (2023) and (Magdy and Jones, 2010b) refer to $\text{R}_{1000}$ as recall-oriented but AP being precision-oriented. In light of the lack of consensus on recall-orientation, in Section 3.1, we propose a new quantitative view of recall-orientation based on the how sensitive a metric is for a user interested in finding every relevant item. This allows us to see recall-orientation along a spectrum and compare the degrees of recall-orientation of different metrics. In Section 3.2, based on this definition, we derive a new recall metric, total search efficiency.

### v6

@@ Abstract.
Although originally developed to evaluate sets of items, recall is often used to evaluate rankings of items, including those produced by recommender, retrieval, and other machine learning systems. The application of recall without a formal evaluative motivation has led to criticism of recall as a vague or inappropriate measure. In light of this debate, we reflect on the measurement of recall in rankings from a formal perspective. Our analysis is composed of three tenets: recall, robustness, and lexicographic evaluation. First, we formally define ‘recall-orientation’ as the sensitivity of a metric to a user interested in finding every relevant item. Second, we analyze recall-orientation from the perspective of robustness with respect to possible content consumers and providers, connecting recall to recent conversations about fair ranking. Finally, we extend this conceptual and theoretical treatment of recall by developing a practical preference-based evaluation method based on lexicographic comparison. Through extensive empirical analysis across multiple recommendation and retrieval tasks, we establish that our new evaluation method, lexirecall, has convergent validity (i.e., it is correlated with existing recall metrics) and exhibits substantially higher sensitivity in terms of discriminative power and stability in the presence of missing labels. Our conceptual, theoretical, and empirical analysis substantially deepens our understanding of recall and motivates its adoption through connections to robustness and fairness.
@@ 1. Introduction
Researchers use ‘recall’ to evaluate rankings across a variety of retrieval, recommendation (Zhao et al., 2022), and machine learning tasks (Sajjadi et al., 2018; Rudinger et al., 2015; Flach and Kull, 2015). ‘Recall at $k$ items’ ($\text{R}_{k}$) and R-Precision (RP) are popular metrics used for measuring recall in rankings. Since the beginning of the ACM Conference on Recommender Systems, on average one third of full papers measure recall in experiments (Figure 1). While there is a colloquial interpretation of recall as measuring coverage (as it might be rightfully interpreted in set retrieval), the research community is far from a principled understanding of recall metrics for rankings. Nevertheless, authors continue to informally refer to evaluation metrics as more or less ‘recall-oriented’ or ‘precision-oriented’ without a formal definition of what this means or quantifying how existing metrics relate to these constructs (Kazai and Lalmas, 2006; Montazeralghaem et al., 2020; Dai et al., 2017; Mohammad et al., 2018; Li et al., 2014; Diaz, 2015).
Figure 1. Percentage of full papers published at the ACM Conference on Recommender Systems that measure recall in experiments. Percentages are macro-averaged across years to control for growth in the number of full papers in the conference proceedings. Details can be found in Appendix A.
@@
Given the prevalence of recall as a metric in recommender system research, understanding recall conceptually, theoretically, and empirically is fundamental to sound evaluation. Indeed, the lack of a principled understanding of or motivation for recall has caused some to question whether recall is useful as a construct at all. Herlocker et al. (2004) suggests that the sparse nature of recommendation data may make recall inappropriate. Cooper (1968) argues that recall-orientation is inappropriate because user search satisfaction depends on the number of items the user is looking for, which may be fewer than all of the relevant items. Zobel et al. (2009) refute several informal justifications for recall: persistence (the depth a user is willing to browse), cardinality (the number of relevant items found), coverage (the number of user intents covered), density (the rank-locality of relevant items), and totality (the retrieval of all relevant items). So, while many ranking experiments compute recall metrics, precisely what and why we are measuring remains unclear.
@@
In this light, we approach the measurement of recall in rankings from a formal perspective, with an objective of proposing a new interpretation of recall with precise conceptual and theoretical grounding. Our analysis is composed of three interrelated tenets: recall, robustness, and lexicographic evaluation. First, we consider recall an essentially contested construct (Gallie, 1955; Jacobs and Wallach, 2021): it is a high level construct believed to be important but with different, conflicting interpretations, as suggested by Zobel et al. (2009). By adopting the interpretation of recall as finding the totality of relevant items, we formally define ‘recall-orientation’ as sensitivity to movement of the bottom-ranked relevant item. Although simple, this definition of recall connects to both early work in position-based evaluation as well as recent work in technology-assisted review. Moreover, by formally defining recall orientation, we can design a new metric, total search efficiency, that precisely measures recall. Second, we consider robustness another essentially contested construct, again with different, conflicting interpretations (Drenkow et al., 2021). By adopting an interpretation of robustness as the effectiveness for the worst-off user, we can connect it to our notion of recall-orientation. We demonstrate that recall—and totality specifically—is aligned with worst-case robustness. Finally, we extend this conceptual and theoretical treatment of recall by developing a practical preference-based evaluation method based on lexicographic comparison. We present a conceptual relationship between recall, robustness, and lexicographic evaluation in Figure 2. Through extensive empirical analysis across various retrieval and recommendation tasks, we establish that our new evaluation method, lexicographic recall or lexirecall, is correlated with existing recall metrics but exhibits substantially higher discriminative power and stability in the presence of missing labels. While conceptually and theoretically grounded in notions of robustness and fairness, lexirecall pragmatically is appropriate when we are concerned with understanding system behavior for users who are focused on finding all relevant items in the same way that experiments adopt reciprocal rank as a high-precision metric. This paper, by focusing on worst-case performance, also complements existing work that has focused on average case (Diaz and Ferraro, 2022) and best-case (Diaz, 2023) preference-based evaluation. Our conceptual, theoretical, and empirical analysis substantially deepens our understanding of recall as a construct and motivates its adoption through connections to robustness and fairness constructs.
@@ 2. Preliminaries
We consider ranking systems that are designed to satisfy users [^1] We adopt user as a general term for searcher in the information retrieval context and content consumer in the multi-stakeholder recommendation context. with with some information need, broadly construed. This information need may be specific or vague, and may or may not be explicit even in the user’s own mind (e.g. “entertain me for the evening” can be considered an information need for home-page recommendations on a video streaming service; “what happened yesterday that I should know about?” for a daily e-mail of recommended new articles from a news publisher). Although users in many recommendation scenarios may not seem to have an information need, they do, at any point in time, have a latent order over items in the catalog. This is true in relatively passive scenarios like radio-like streaming audio where a user, although not explicitly expressing or thinking about their preferences, nevertheless reveals them through their consumption behavior.
While a user’s information need is never directly revealed to the ranking system, the user expresses an observable request to the ranking system. A request can include information provided by the user either explicitly (e.g., a query or question for text-based search; an application or site context for recommendation) or implicitly (e.g., geo-location for text-based search; engagement or session history for recommendation (Ludewig and Jannach, 2018)) or both (Zamani and Croft, 2020; Sontag et al., 2012).
@@
The objective of ranking evaluation is to determine the quality of a permutation $\pi\in S_{n}$ for the user. In the remainder of this section, we will detail precisely how we do this. [^2] While ranking does not cover all recommendation scenarios, it covers many, including personalized search, home-page recommendations, related product recommendations, and social media timeline ranking, and is frequently considered in research; further, non-ranking interfaces such as music or video streams may well be implemented as a final selection or transformation of an underlying (partial) ranking, and therefore better understanding of ranking evaluation may be applicable even to recommendation surfaces that do not directly expose the ranking to the user.
@@ 2.1. Relevance
The relevance of an item refers to its value with respect to a user’s information need. In this work, we focus on binary relevance, an approach regularly used in information retrieval and recommender system literature, especially when measuring recall. [^3] We discuss how our analysis extends in ordinal grades and preferences in Section 7.2.3. Let $\mathcal{R}\subset\mathcal{D}$ be the set of items labeled relevant to the request where $m=|\mathcal{R}|$.
@@
Figure 3. Relevance Projection with Imputation. Given a set of relevant item ids =$\mathcal{R}=\{083,107,511,668,820,977\}$, relevance projection of an incomplete top-10 ranking $\pi$ and relevant set $\mathcal{R}$ to a $m\times 1$ vector of positions $p$. We also show the inverse projection vector $\overline{p}$ of items at specific recall levels.
@@ 2.2. Permutation Imputation
Pessimistic imputation is a conservative placement of the unretrieved relevant items and is well-aligned with our interest in robust performance. Moreover, it is consistent with behavior of metrics like rank-biased precision (Moffat and Zobel, 2008), which implicitly applies an exposure of 0 for unretrieved relevant items (i.e., for large $n$, $\lim_{i\rightarrow n}\gamma^{i}\approx 0$); and average precision [^4] As defined in trec_eval., which implicitly applies an exposure of 0 for unretrieved relevant items (i.e., for large $n$, $\lim_{i\rightarrow n}\frac{1}{i}\approx 0$). We show an example of projection with imputation in Figure 3.
@@ 2.3. Measuring Effectiveness
When evaluating a ranking system, we consider two sets of important stakeholders: users and providers. Users approach the system with information needs and requests and ultimately define what is relevant. Providers contribute items to the search system which serve to satisfy users. Each item in the corpus is attributable, explicitly or not, to a content provider. We include providers in our treatment to draw connections in existing work in the fair ranking literature, which often focuses on providers as stakeholders (Ekstrand et al., 2022). In this section, we will characterize a broad family of evaluation metrics for these two sets of users that will allow us, in subsequent sections, to define formal notions of recall and robustness.
@@ 2.3.1. Measuring Effectiveness for Users
The product $e(p_{i})z(i,m)$ is a decomposition of what Carterette refers to as a ‘discount function’ into an explicit function that models exposure and another that addresses any recall normalization (Carterette, 2011). Within the set of recall-level metrics, we are further interested in the sub-class of metrics that satisfy the following criteria for ‘top-heaviness’.
Definition 2.2.
@@
Top-heaviness indicates that, in the event that there are $j$ unjudged, relevant items in positions above the remaining $m-j$ relevant items, the metric computed over all $m$ items must be greater than or equal to the metric value computed over only the $m-j$ items. Because we deal with metrics that include functions of $m$, this is not an obvious property, but one that will be important as we consider incomplete judgments and relationships between possible users in Section 4.555For more information on the relationship between incomplete judgments and metric stability, see
@@
Beyond classic ranking metrics, top-heavy recall-level metrics include non-traditional metrics such as those based on linear discounting (e.g., $e_{\text{lin}}(i)=1-\frac{i}{n}$). This results in a much broader class of metrics than those normally considered, for example, in the formal analysis of ranking metrics (Moffat, 2013; Amigó et al., 2013; Ferrante et al., 2015). As a result, while all top-heavy recall-level metrics satisfy some formal properties of ranking evaluation metrics, large subsets of top-heavy recall-level metrics may satisfy more. A detailed analysis of formal properties of top-heavy recall-level metrics can be found in Appendix B. As mentioned before, we can contrast this with Carterette’s decomposition which focuses on the decomposition of metrics into gain and discount components (Carterette, 2011). In our case, we do not model gain, since we deal with binary relevance. Our exposure and normalization functions, then, precisely define a subset of Carterette’s discount functions that do not fit into his metric taxonomy since they do not consider recall normalization (Carterette, 2011).
@@ 2.3.2. Measuring Effectiveness for Providers
where $m$ and $p$ are based on $\mathcal{R}$. Unless necessary, we will drop the subscript $e$ from $\eta$ for clarity.
@@ 2.4. Evaluation Method Desiderata
    • System sensitivity. Is the evaluation method better able to distinguish between rankers compared to existing methods? (Section 6.2.3)
@@ 3. Recall
As mentioned in Section 1, the description of ranked ranking metrics as ‘recall-oriented’ remains poorly defined, leaving the formal analysis of metrics for recall-orientation difficult. From a technical point of view, some work considers recall-orientation to be a binary criteria, dependent on whether a metric includes the recall base (i.e., $\mathcal{R}$) in order to be computed (Kazai and Lalmas, 2006; Sakai, 2006). This would include metrics that compute set-based recall at some rank cutoff (Tomlinson and Hedin, 2017; Cormack and Grossman, 2017) as well as metrics like AP and $\mathrm{NDCG}$. Using a set-based recall metric is particularly well-suited for recall-orientation in early stages of multi-stage ranking (Macdonald et al., 2013; Mohammad et al., 2018). A binary notion of recall-orientation does not capture that some metrics may be more recall-oriented than others. This is captured, in part, by references to recall-orientation as related to the depth in the ranking considered by the user (Diaz, 2015). More frequently, authors appeal to metrics like AP and $\text{R}_{1000}$ as being recall-oriented without clear discussion of what this means (Montazeralghaem et al., 2020; Dai et al., 2017; Li et al., 2014; Hashemi et al., 2020). On the other hand, both Mackie et al. (2023) and (Magdy and Jones, 2010b) refer to $\text{R}_{1000}$ as recall-oriented but AP being precision-oriented. In light of the lack of consensus on recall-orientation, in Section 3.1, we propose a new quantitative view of recall-orientation based on how sensitive a metric is for a user interested in finding every relevant item. This allows us to see recall-orientation along a spectrum and compare the degrees of recall-orientation of different metrics. In Section 3.2, based on this definition, we derive a new recall metric, total search efficiency.